OpenStudy (anonymous):

solve the equation 10t^2=2r^3-72t

OpenStudy (solomonzelman):

Do you mean this: \(\large\color{black}{ \displaystyle 10t^2=2\color{red}{t}^3-72t }\) ? (you said r instead of t)

OpenStudy (solomonzelman):

Or, is it the way you wrote it: \(\large\color{black}{ \displaystyle 10t^2=2r^3-72t }\) And you want to solve for \(r\), or for \(t\)?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

the first thing I said?